### v2

@@ 3.3 Exact reconstruction of the intrusive ROM operators
To show that our novel approach exactly reconstructs the intrusive ROM operators, we first show that the FOM snapshot data generated in our proposed approach is equivalent to snapshot data generated from the intrusive ROM itself, which constitutes the ideal inference data.
For this purpose, we insert the definition of the rank-ensuring initial conditions $\bar{\bm{x}}_{s}$ (17) and the explicit Euler FOM time step (18) into the definition of the explicit Euler time derivative snapshot (19) to find
@@
Since ${\mathbf{P}}$ has full rank, the solution $\hat{\mathbf{O}}=\tilde{\mathbf{O}}$ is unique, so the method exactly reconstructs the intrusive ROM operator.
@@ 3.4 Assumptions on the FOM solver
Operator inference is a nonintrusive model reduction concept developed for FOM solvers that do not provide the access to FOM operators required by intrusive model reduction. Consequently, imposing strong assumptions on the FOM solver contradicts the core motivation for operator inference. In our proposed approach, we only impose two assumptions. First, we assume that specific initial conditions can be imposed; concretely, linear combinations of the ROM basis modes (17). Second, we assume that explicit Euler can be chosen as time discretization method to generate the operator inference snapshot data. For other explicit time discretizations, it is unclear whether a consistent time derivative snapshot computation is possible. For implicit methods, on the other hand, the full rank of ${\mathbf{P}}$ might not be guaranteed. For example, if the FOM time steps (18) are replaced by implicit Euler steps,
@@
then the matrix ${\mathbf{P}}$ should, for time discretization consistency, be composed of vectors ${\bm{p}}({\bm{x}}_{1}^{s},{\bm{u}}_{1}^{s})$ instead of ${\bm{p}}({\bm{x}}_{0}^{s},{\bm{u}}_{0}^{s})$. However, in contrast to the initial conditions ${\bm{x}}_{0}^{s}$, we do not have full control over the states ${\bm{x}}_{1}^{s}$ after one time step. Consequently, we generally cannot guarantee the full rank of ${\mathbf{P}}$.
Note that this requirement on the time discretization only affects the generation of operator inference snapshot data, but neither the generation of POD snapshot data, nor the simulations that are performed with the resulting ROM.
Compared with the re-projection approach [18], our proposed approach is less intrusive because it does not require a re-projection step after each time step.
@@ 3.6 Choice of time step size $\Delta t$
Since we only perform a single time step per trajectory, we do not need to take stability into account when determining the time step size $\Delta t$. In exact arithmetics, the size of $\Delta t$ also does not affect the accuracy of the inferred operators since the effect of $\Delta t$ in computing ${\bm{x}}_{1}^{s}$ in (18) cancels out when computing the time derivative snapshot $\dot{\breve{\bm{x}}}_{1}^{s}$ in (19). In floating point arithmetics, however, the accuracy of the inferred operators depends on how accurate the intermediate results can be represented as floating-point numbers. In particular, if $\|\tilde{\mathbf{O}}\|_{2}$ is very large or very small, the differences $\breve{\bm{x}}_{1}^{s}-\breve{\bm{x}}_{0}^{s}$ cannot be represented well as floating-point numbers. To counteract this effect, $\Delta t=\frac{1}{\|\tilde{\mathbf{O}}\|_{2}}$ would be a good choice. Since we do not know $\tilde{\mathbf{O}}$ a priori, we can only approximate $\|\tilde{\mathbf{O}}\|_{2}$ based on the POD snapshot data.
@@
for ROM dimension $n=1$.
@@ 4 Numerical experiments
The following numerical experiments demonstrate that Algorithm 1 exactly reconstructs the corresponding intrusive ROM operators. Many existing works on operator inference assess the accuracy of the inferred ROMs based on trajectory errors. As shown in [24], trajectory errors can be misleading because they are sensitive to the choice of trajectories. Furthermore, the objective of operator inference is to reconstruct the intrusive ROM, so this intrusive ROM should serve as ground truth, not the underlying FOM. We therefore assess the reconstruction accuracy with relative operator errors $\frac{\|\hat{\mathbf{O}}-\tilde{\mathbf{O}}\|_{F}}{\|\tilde{\mathbf{O}}\|_{F}}$.
@@ 4.1 Chafee-Infante equation / Allen-Cahn equation
(28) $x(0,t)=u(t),\hskip 10.00002pt\frac{\partial}{\partial t}x(1,t)=0,\hskip 10.00002pt\hskip 10.00002ptt\in[0,0.1],$
@@
This system is polynomial of degree $l=3$ with ${\mathbf{A}}_{1}\in\mathbb{R}^{N\times N}$, ${\mathbf{A}}_{2}\in\mathbb{R}^{N\times N_{2}}$ with $N_{2}=\frac{N(N+1)}{2}$, ${\mathbf{A}}_{3}\in\mathbb{R}^{N\times N_{3}}$ with $N_{3}=\frac{N(N+1)(N+2)}{6}$, ${\mathbf{B}}\in\mathbb{R}^{N\times{N_{u}}}$ with ${N_{u}}=1$, the input signal $u_{k}=10(\sin(k\pi\Delta t)+1)$, $k=0,\dots,K-1$ and $K=10^{4}$. The resulting $110^{4}+$ 1 snapshots are used to compute the POD bases ${\mathbf{V}}_{n}\in\mathbb{R}^{N\times n}$ with $n=1,\dots,14$.
We apply Algorithm 1 with $\Delta t=3.2733\cdot 10^{-5}$ according to estimate (24) and compute the relative operator errors for ROM dimensions $n=1,\dots,14$. Fig. 2 shows that the operator error is on the order of machine precision and Fig. 2 shows that the data matrices ${\mathbf{P}}$ have moderate condition numbers.
@@ 4.2 Ice sheet model
(30) $\frac{\partial}{\partial t}x=c_{1}x^{2}\partial_{\xi}x+c_{2}x^{5}|\partial_{\xi}x|^{2}\partial_{\xi}x$
@@
We apply Algorithm 1 with $\Delta t=1.4726\cdot 10^{13}$ according to estimate (24). This extreme value can be explained by the extreme operator norm $\|\tilde{\mathbf{O}}\|_{2}=4.271\cdot 10^{-14}$ for ROM dimension $n=1$. The relative operator errors for ROM dimensions $n=1,\dots,7$ are depicted in Fig. 2. Like in the previous test case, these operator errors are essentially on the order of machine precision. The fact that these errors are a few orders of magnitude larger than the errors in the previous test case is presumably related to the larger condition numbers as depicted in Fig. 2.
For ROM dimension $n=7$, the proposed method performs ${n_{f}}=3087$ FOM time steps. This number is larger than the $2001$ FOM time steps with re-projection reported in [2]. However, the authors also report that this budget of snapshots is depleted before the inference of the ROM operators of dimension $n=6$ is completed. Hence, the ROM operators of dimensions $n=6$ and $n=7$ are partially computed with regularized operator inference and snapshot data polluted with non-Markovian terms. Consequently, the inferred operators of dimension $n=6$ and $n=7$ are not reconstructed exactly. The operators of dimension $n\leq 5$ might be reconstructed exactly, but the used $2001$ snapshots are more than the ${n_{f}}=365$ snapshots used by our method for dimension $n=5$.
@@ 4.3 Burgers’ equation
with state $x\in\mathbb{R}$, spatial coordinate $\xi\in\Omega=(-1,1)$, time $t\in[0,1]$. We impose periodic boundary conditions and the initial condition $x(\xi,0)=-\sin\left(\frac{\pi}{2}\xi\right)$ for all $\xi\in\Omega$.
@@
This system is polynomial of degree $l=2$ with ${\mathbf{A}}_{1}\in\mathbb{R}^{N\times N}$, ${\mathbf{A}}_{2}\in\mathbb{R}^{N\times N_{2}}$ with $N_{2}=\frac{N(N+1)}{2}$ and $K=10^{4}$. Concretely, we use a spatial discretization as described in [28] such that the diffusion matrix ${\mathbf{A}}_{1}$ is symmetric and negative semi-definite, and the convection operator ${\mathbf{A}}_{2}$ satisfies the skew-symmetry property ${\bm{x}}^{T}{\mathbf{A}}_{2}{\bm{x}}^{2}=0$ for all ${\bm{x}}\in\mathbb{R}^{N}$. The resulting $110^{4}+$ 1 snapshots are used to compute the POD bases ${\mathbf{V}}_{n}\in\mathbb{R}^{N\times n}$ with $n=1,\dots,10$.
We perform Algorithm 1 with $\Delta t=0.1013$ according to estimate (24) and compute the relative operator errors for ROM dimensions $n=1,\dots,10$. Fig. 2 shows operator errors on the order of machine precison and Fig. 2 very small condition numbers.
In addition, we observe that both, the intrusive and the inferred operators preserve the skew-symmetry of the FOM convection operator and the symmetry and negative semi-definiteness of the FOM diffusion operator. Fig. 3 shows the energy-preserving constraint violation $\sum_{1\leq i,j,k\leq n}\big{|}\hat{h}_{ijk}+\hat{h}_{jik}+\hat{h}_{kji}\big{|}$ as defined in [12] for the inferred operator with $\hat{h}_{ijk}={\bm{e}}_{i}\hat{\mathbf{A}}_{2}{\mathbf{I}}_{N}^{(2)}({\bm{e}}_{j}\otimes{\bm{e}}_{k})$ and for the intrusive operator with $\hat{h}_{ijk}={\bm{e}}_{i}\tilde{\mathbf{A}}_{2}{\mathbf{I}}_{N}^{(2)}({\bm{e}}_{j}\otimes{\bm{e}}_{k})$. As can be seen, this violation is essentially on the order of machine precision for both, inferred and intrusive operators. Fig. 5 shows the symmetry violation of the inferred and the intrusive diffusion matrix, $\frac{\left\|\hat{\mathbf{A}}_{1}-\hat{\mathbf{A}}_{1}^{T}\right\|}{\left\|\hat{\mathbf{A}}_{1}\right\|}$ and $\frac{\left\|\tilde{\mathbf{A}}_{1}-\tilde{\mathbf{A}}_{1}^{T}\right\|}{\left\|\tilde{\mathbf{A}}_{1}\right\|}$, respectively. Also this violation is on the order of machine precision. Consequently, the diffusion matrices are symmetric, so their eigenvalues are real. These eigenvalues - multiplied with $-1$ are depicted in Fig. 5. As can be seen the eigenvalues of the inferred diffusion matrices coincide with the eigenvalues of the corresponding intrusive matrices. Hence, both ROMs in particular preserve the negative semi-definiteness of the FOM diffusion operator.
Figure 2: Condition numbers of data matrices ${\mathbf{P}}$.
Figure 5: Eigenvalues of diffusion matrix of Burgers’ equation multiplied with $-1$.

### v3

@@ 3.3 Exact reconstruction of the intrusive ROM operators
By combining the rank-ensuring initial conditions and input signals in Section 3.1 and the single-step ensemble in Section 3.2, our novel approach exactly reconstructs the intrusive ROM operators, as stated by the following theorem.
[…]
For any $n\in\mathbb{N}^{+}$, ${N_{u}}\in\mathbb{N}_{0}$, $\mathcal{I}\subset\mathbb{N}_{0}$ and $\Delta t>0$, the operator inference minimization problem (8) with snapshot data $(\breve{\bm{x}}_{k},\dot{\breve{\bm{x}}}_{k},{\bm{u}}_{k})=(\bar{\bm{x}}_{s},\dot{\bar{\bm{x}}}_{s},\bar{\bm{u}}_{s})$ defined in (14) and (19) has a unique solution. This solution is equivalent to the intrusive ROM operators defined in (4) and (5).
Proof.
We first show that the FOM snapshot data generated in our proposed approach is equivalent to snapshot data generated from the intrusive ROM itself, which constitutes the ideal inference data. For this purpose, we insert the definition of the rank-ensuring initial conditions $\bar{\bm{x}}_{s}$ (17) and the explicit Euler FOM time step (18) into the definition of the explicit Euler time derivative snapshot (19) to find
@@
Since $\mathbf{P}$} has full rank as shown in Theorem 3.1, the solution $\hat{\mathbf{O}}=\tilde{\mathbf{O}}$ is unique, so the method exactly reconstructs the intrusive ROM operator.
@@ 3.4 Assumptions on the FOM solver
Operator inference is a nonintrusive model reduction concept developed for FOM solvers that do not provide the access to FOM operators required by intrusive model reduction. Consequently, imposing strong assumptions on the FOM solver contradicts the core motivation for operator inference. In our proposed approach, we only impose two assumptions.
First, we assume that specific initial conditions can be imposed; concretely, linear combinations of the ROM basis modes (17). For POD bases, these modes are linear combinations of FOM snapshots. Since these snapshots are physically realizable, the specific initial conditions are physically realizable as well in most cases. Exceptions include models that only allow a certain range of state values, for example positivity of temperature, or require specific boundary conditions. These exceptions are beyond the scope of this article since they typically require some additional treatment already on the intrusive ROM level, for example the introduction of a lifting function [22] to satisfy boundary conditions.
Second, we assume that explicit Euler can be chosen as time discretization method to generate the operator inference snapshot data. However, this restriction can be relaxed to higher-order Runge-Kutta schemes if input signals are interpolated between time steps. These schemes can be written as explicit Euler schemes whose right-hand side has higher polynomial degree [18]. Consequently, our approach can be applied to explicit Runge-Kutta methods (and multi-step methods that use an explicit Runge-Kutta method to compute the first time step). For implicit methods, on the other hand, the full rank of $\mathbf{P}$} might not be guaranteed. For example, if the FOM time steps (18) are replaced by implicit Euler steps,
@@
then the matrix $\mathbf{P}$} should, for time discretization consistency, be composed of vectors ${\bm{p}}({\bm{x}}_{1}^{s},{\bm{u}}_{1}^{s})$ instead of ${\bm{p}}({\bm{x}}_{0}^{s},{\bm{u}}_{0}^{s})$. However, in contrast to the initial conditions ${\bm{x}}_{0}^{s}$, we do not have full control over the states ${\bm{x}}_{1}^{s}$ after one time step. Consequently, we generally cannot guarantee the full rank of $\mathbf{P}$}. Alternatively, we can construct the matrix $\mathbf{P}$} from the vectors ${\bm{p}}({\bm{x}}_{0}^{s},{\bm{u}}_{0}^{s})$. This way, we guarantee the full rank of $\mathbf{P}$} but introduce a time discretization inconsistency that impedes exact reconstruction of the intrusive ROM operators.
Note that this requirement on the time discretization only affects the generation of operator inference snapshot data, but neither the generation of POD snapshot data, nor the simulations that are performed with the resulting ROM. The POD snapshot data can be generated from any FOM time discretization, and - like intrusive ROM operators - the inferred operators can be used for any ROM time discretization.
Note further that the time step size $\Delta t$ is provided as an input in Algorithm 1 only for ease of presentation. Our proposed approach is also applicable to FOM solver that do not allow the users to specify the time step size themselves. In such cases, $\Delta t$ is not provided as an input and the end time $T$ cannot be chosen to be equal to $\Delta t$. Consequently, we can generally not guarantee that the FOM solver only performs a single time step, but by setting $T$ to a small value, we can limit the number of excess time steps. From the solver output, we then only use the state snapshot ${\bm{x}}_{1}$ after the first time step and the corresponding time $t_{1}$ to compute the time step size $\Delta t$.
While our approach differs decisively from the re-projection approach [18] in terms of control over which snapshots are generated, both approaches are very similar in terms of intrusiveness. In fact, our approach can be seen as the re-projection approach performed for $n_{f}$} separate trajectories, but stopped after a single time step and before the first re-projection step. Consequently, the described assumptions on the FOM solver are almost the same as for the re-projection approach. The only exception are implicit time discretization schemes for linear models which are allowed in re-projection but not in our approach because they inhibit control over the generated snapshots.
@@ 3.6 Choice of time step size $\Delta t$
As stated in Theorem 3.3, our proposed method achieves exact reconstruction for any $\Delta t>0$. Indeed, in exact arithmetics, the size of $\Delta t$ also does not affect the accuracy of the inferred operators since the effect of $\Delta t$ in computing ${\bm{x}}_{1}^{s}$ in (18) cancels out when computing the time derivative snapshot $\dot{\breve{\bm{x}}}_{1}^{s}$ in (19). In floating point arithmetics, however, the accuracy of the inferred operators depends on how accurate the intermediate results can be represented as floating-point numbers. If $\Delta t$ is determined by the FOM solver, we have to accept the inaccuracy in the ROM operator reconstruction induced by this limitations of floating-point number representation. However, if $\Delta t$ can be specified by the user, we can choose $\Delta t$ such that these limitations are mitigated.
Concretely, the differences $\breve{\bm{x}}_{1}^{s}-\breve{\bm{x}}_{0}^{s}$ cannot be represented well as floating-point numbers if $\|\tilde{\mathbf{O}}\|_{2}$ is very large or very small. To counteract this effect, $\Delta t=\frac{1}{\|\tilde{\mathbf{O}}\|_{2}}$ would be a good choice. Since we do not know $\tilde{\mathbf{O}}$ a priori, we can only approximate $\|\tilde{\mathbf{O}}\|_{2}$ based on the POD snapshot data.
@@
for ROM dimension $n=1$, so ${\mathbf{V}}_{n}={\bm{v}}_{1}$.
Note that we do not need to take stability into account when determining the time step size $\Delta t$ since we only perform a single time step per trajectory.
@@ 4 Numerical experiments
The following numerical experiments demonstrate that Algorithm 1 exactly reconstructs the corresponding intrusive ROM operators. Many existing works on operator inference assess the accuracy of the inferred ROMs based on trajectory errors. As shown in [25], trajectory errors can be misleading because they are sensitive to the choice of trajectories. Furthermore, the objective of operator inference is to reconstruct the intrusive ROM, so this intrusive ROM should serve as ground truth, not the underlying FOM (see Appendix C). We therefore assess the reconstruction accuracy with relative operator errors $\frac{\|\hat{\mathbf{O}}-\tilde{\mathbf{O}}\|_{F}}{\|\tilde{\mathbf{O}}\|_{F}}$.
@@
Figure 2: Condition numbers of data matrices $\mathbf{P}$}.
@@ 4.1 Chafee-Infante equation / Allen-Cahn equation
(28) $x(0,t)=u(t),\hskip 10.00002pt\frac{\partial}{\partial{\xi}}x(1,t)=0,\hskip 10.00002pt\hskip 10.00002ptt\in[0,0.1],$
@@
This system is polynomial of degree $l=3$ with ${\mathbf{A}}_{1}\in\mathbb{R}^{N\times N}$, ${\mathbf{A}}_{2}\in\mathbb{R}^{N\times N_{2}}$ with $N_{2}=\frac{N(N+1)}{2}$, ${\mathbf{A}}_{3}\in\mathbb{R}^{N\times N_{3}}$ with $N_{3}=\frac{N(N+1)(N+2)}{6}$, ${\mathbf{B}}\in\mathbb{R}^{N\times{N_{u}}}$ with ${N_{u}}=1$, the input signal $u_{k}=10(\sin(k\pi\Delta t)+1)$, $k=0,\dots,K-1$ and $K=10^{4}$. The resulting $10^{4}+1$ snapshots are used to compute the POD bases ${\mathbf{V}}_{n}\in\mathbb{R}^{N\times n}$ with $n=1,\dots,14$.
We apply Algorithm 1 with $\Delta t={3.2741}\cdot 10^{-5}$ according to estimate (24) and compute the relative operator errors for ROM dimensions $n=1,\dots,14$. Fig. 2 shows that the operator error is on the order of machine precision and Fig. 2 shows that the data matrices $\mathbf{P}$} have moderate condition numbers.
@@ 4.2 Ice sheet model
(30) ${\frac{\partial}{\partial t}x(\xi,t)=c_{1}{\frac{\partial}{\partial\xi}\Bigg(}x^{2}(\xi,t)\frac{\partial}{\partial\xi}x(\xi,t){\Bigg)}+c_{2}{\frac{\partial}{\partial\xi}\Bigg(}x^{5}(\xi,t)\left|\frac{\partial}{\partial\xi}x(\xi,t)\right|^{2}\frac{\partial}{\partial\xi}x(\xi,t){\Bigg)}}$
@@
We apply Algorithm 1 with $\Delta t={1.4773}\cdot 10^{13}$ according to estimate (24). This extreme value can be explained by the extreme operator norm $\|\tilde{\mathbf{O}}\|_{2}=4.271\cdot 10^{-14}$ for ROM dimension $n=1$. The relative operator errors for ROM dimensions $n=1,\dots,7$ are depicted in Fig. 2. Like in the previous test case, these operator errors are essentially on the order of machine precision. The fact that these errors are a few orders of magnitude larger than the errors in the previous test case is presumably related to the larger condition numbers as depicted in Fig. 2.
For ROM dimension $n=7$, the proposed method performs ${n_{f}}=3087$ FOM time steps. This number is larger than the $2001$ FOM time steps with re-projection reported in [2]. However, the authors also report that this budget of snapshots is depleted before the inference of the ROM operators of dimension $n=6$ is completed. Hence, their ROM operators of dimensions $n=6$ and $n=7$ are partially computed with regularized operator inference and snapshot data polluted with non-Markovian terms. Consequently, their inferred operators of dimension $n=6$ and $n=7$ are not reconstructed exactly. Their operators of dimension $n\leq 5$ might be reconstructed exactly, but the used $2001$ snapshots are more than the ${n_{f}}=365$ snapshots used by our method for dimension $n=5$.
@@ 4.3 Burgers’ equation
Figure 5: Eigenvalues of diffusion matrix of Burgers’ equation multiplied with 1-1.
@@
with spatial coordinate $\xi\in\Omega=(-1,1)$, time $t\in[0,1]$. We impose periodic boundary conditions and the initial condition $x(\xi,0)=-\sin\left({\pi}\xi\right)$ for all $\xi\in\Omega$.
@@
This system is polynomial of degree $l=2$ with ${\mathbf{A}}_{1}\in\mathbb{R}^{N\times N}$, ${\mathbf{A}}_{2}\in\mathbb{R}^{N\times N_{2}}$ with $N_{2}=\frac{N(N+1)}{2}$ and $K=10^{4}$. Concretely, we use a spatial discretization as described in [30] such that the diffusion matrix ${\mathbf{A}}_{1}$ is symmetric and negative semi-definite, and the convection operator ${\mathbf{A}}_{2}$ satisfies the skew-symmetry property ${\bm{x}}^{T}{\mathbf{A}}_{2}{\bm{x}}^{2}=0$ for all ${\bm{x}}\in\mathbb{R}^{N}$. The resulting $10^{4}+1$ snapshots are used to compute the POD bases ${\mathbf{V}}_{n}\in\mathbb{R}^{N\times n}$ with $n=1,\dots,10$.
We perform Algorithm 1 with $\Delta t={1.0134\cdot 10^{-1}}$ according to estimate (24) and compute the relative operator errors for ROM dimensions $n=1,\dots,10$. Fig. 2 shows operator errors on the order of machine precison and Fig. 2 very small condition numbers. In addition, we observe that both, the intrusive and the inferred operators preserve the skew-symmetry of the FOM convection operator and the symmetry and negative semi-definiteness of the FOM diffusion operator. Fig. 3 shows the energy-preserving constraint violation $\sum_{1\leq i,j,k\leq n}\big|\hat{h}_{ijk}+\hat{h}_{jik}+\hat{h}_{kji}\big|$ as defined in [12] for the inferred operator with $\hat{h}_{ijk}={\bm{e}}_{i}\hat{\mathbf{A}}_{2}{\mathbf{I}}_{N}^{(2)}({\bm{e}}_{j}\otimes{\bm{e}}_{k})$ and for the intrusive operator with $\hat{h}_{ijk}={\bm{e}}_{i}\tilde{\mathbf{A}}_{2}{\mathbf{I}}_{N}^{(2)}({\bm{e}}_{j}\otimes{\bm{e}}_{k})$. As can be seen, this violation is essentially on the order of machine precision for both, inferred and intrusive operators. Fig. 5 shows the symmetry violation of the inferred and the intrusive diffusion matrix, $\frac{\left\|\hat{\mathbf{A}}_{1}-\hat{\mathbf{A}}_{1}^{T}\right\|}{\left\|\hat{\mathbf{A}}_{1}\right\|}$ and $\frac{\left\|\tilde{\mathbf{A}}_{1}-\tilde{\mathbf{A}}_{1}^{T}\right\|}{\left\|\tilde{\mathbf{A}}_{1}\right\|}$, respectively. Also this violation is on the order of machine precision. Consequently, the diffusion matrices are symmetric, so their eigenvalues are real. These eigenvalues - multiplied with 1-1 - are depicted in Fig. 5. As can be seen the eigenvalues of the inferred diffusion matrices coincide with the eigenvalues of the corresponding intrusive matrices. Hence, both ROMs in particular preserve the negative semi-definiteness of the FOM diffusion operator.